OpenStudy (anonymous):

The formula for the circumference C of a circle with radius r is C = 2πr. What is the formula for finding π (pi) in terms of the circumference and radius?

OpenStudy (anonymous):

π = C + 2r π = C - 2r π = two r divided by C π = C divided by two r